# Benchmarking Embedding Aggregation Methods in Computational Pathology: A Clinical Data Perspective

### v1

@@ Abstract
Recent advances in artificial intelligence (AI), in particular self-supervised learning of foundation models (FMs), are revolutionizing medical imaging and computational pathology (CPath). A constant challenge in the analysis of digital Whole Slide Images (WSIs) is the problem of aggregating tens of thousands of tile-level image embeddings to a slide-level representation. Due to the prevalent use of datasets created for genomic research, such as TCGA, for method development, the performance of these techniques on diagnostic slides from clinical practice has been inadequately explored. This study conducts a thorough benchmarking analysis of ten slide-level aggregation techniques across nine clinically relevant tasks, including diagnostic assessment, biomarker classification, and outcome prediction. The results yield following key insights: (1) Embeddings derived from domain-specific (histological images) FMs outperform those from generic ImageNet-based models across aggregation methods. (2) Spatial-aware aggregators enhance the performance significantly when using ImageNet pre-trained models but not when using FMs. (3) No single model excels in all tasks and spatially-aware models do not show general superiority as it would be expected. These findings underscore the need for more adaptable and universally applicable aggregation techniques, guiding future research towards tools that better meet the evolving needs of clinical-AI in pathology. The code used in this work are available at https://github.com/fuchs-lab-public/CPath_SABenchmark.
@@ 2 Methods
To select aggregation methods for our study, we conducted a literature search on Google Scholar with the query “((deep learning) AND ((computational histopathology) OR (whole slide images)) AND (classification))” for publications between 2017 and 2023. While recognizing the potential of hierarchical models to fuse local and global features, we focused on embedding aggregation technologies that do not require instance-level labels, drawing from studies that utilize embeddings at a uniform magnification for consistency. Our analysis centers on methods with public, implementable code and those frequently used in recent scholarly work for benchmarking purposes. Details on the selected aggregation methods and their computational complexities are provided in Appendix 2. Table 3.
@@ 3 Summary of Datasets
1. BCa: Breast cancer (BCa) detection cohort. Breast cancer blocks and normal breast blocks were obtained from the pathology laboratory information system. A total of 1998 slides were sampled, with 999 positive and 999 negative. The positive slides were selected from blocks that received the routine biomarker panel for cancer cases (estrogen receptor: ER, progesterone receptor: PR, human epidermal growth factor receptor 2: HER2, and Ki67), while negative slides were selected from breast cases that did not have an order for the routine panel. Additionally, negative cases were selected if they were not mastectomy cases, did not have a synoptic report associated with the case, and had no mention of cancer or carcinoma in the report. 2. IBD: Inflammatory Bowel Disease (IBD) detection cohort. Normal mucosa samples were obtained from patients undergoing screening and routine surveillance lower endoscopy from 2018 to 2022. IBD cases, including first diagnoses and follow-ups, were included. A total of 1441 slides were sampled, 717 with active inflammation and 724 with normal mucosa. 3. BCa ER, BCa PR, BCa HER2: BCa biomarker prediction cohorts. Breast cancer cases are routinely assessed for ER, PR, and HER2 status using immunohistochemistry (IHC) and Fluorescence In Situ Hybridization (FISH). Results for each biomarker were automatically extracted from the pathology report. 4. BioMe BR HRD:: Breast (BR) Homologous Repair Deficiency (HRD) prediction cohort. Mount Sinai BioMe is a whole-exome sequencing cohort of 30k individuals, where carriers of pathogenic and protein-truncating variants affecting HRD genes, i.e., BRCA1 BRCA2 BRIP1 PALB2 RAD51 RAD51C RAD51D ATM ATR CHEK1 CHEK2, where included as positives. A subset of the BioMe dataset of patients with available breast pathology slides were included. Slides containing solely normal breast tissue and slides with breast cancer were both included. 5. LUAD EGFR: EGFR (Epidermal Growth Factor Receptor) mutation status prediction in Lung Adenocarcinoma (LUAD). Two datasets were collected, one from MSHS and one from MSKCC. For the MSHS dataset, a total of 294 slides were obtained from MSHS clinical slide database, 103 positive and 191 negative. The cohort was built following the guidelines described in previous work [31] to map mutations to a binary target. The MSKCC dataset consists of 1,000 slides with 307 positive and 693 negative and is a random subset of the dataset described in [3]. See [3] for additional details. 6. NSCLC IO: Lung cancer immunotherapy response prediction. Non-small cell lung cancer (NSCLC) patients who received PD-L1 blockade-based immunotherapy were considered. Cytology specimens were excluded. The objective overall response was determined by RECIST [32] and performed by a blinded thoracic radiologist. A total of 454 slides were obtained, 86 positive and 368 negative.
@@ Acknowledgements
This work is supported in part through the use of research platform AI-Ready Mount Sinai (AIR.MS) and the expertise provided by the team at the Hasso Platner Institute for Digital Health at Mount Sinai (HPI.MS). This work was supported in part through the computational and data resources and staff expertise provided by Scientific Computing and Data at the Icahn School of Medicine at Mount Sinai and supported by the Clinical and Translational Science Awards (CTSA) grant UL1TR004419 from the National Center for Advancing Translational Sciences.

### v2

@@ Abstract
Recent advances in artificial intelligence (AI), in particular self-supervised learning of foundation models (FMs), are revolutionizing medical imaging and computational pathology (CPath). A constant challenge in the analysis of digital Whole Slide Images (WSIs) is the problem of aggregating tens of thousands of tile-level image embeddings to a slide-level representation. Due to the prevalent use of datasets created for genomic research, such as TCGA, for method development, the performance of these techniques on diagnostic slides from clinical practice has been inadequately explored. This study conducts a thorough benchmarking analysis of nine slide-level aggregation techniques across six clinically relevant tasks, including diagnostic assessment, biomarker classification, and outcome prediction. The results yield following key insights: (1) Embeddings derived from domain-specific (histological images) foundation models outperform those from generic ImageNet-based models across aggregation methods. (2) Spatial-aware aggregators enhance the performance significantly when using ImageNet pre-trained models. (3) No single model excels in all tasks, though spatial-aware models show general superiority. These findings underscore the need for more adaptable and universally applicable aggregation techniques, guiding future research towards tools that better meet the evolving needs of clinical-AI in pathology. The code used in this work are available at https://github.com/fuchs-lab-public/CPath_SABenchmark.
@@ 2 Methods
To select aggregation methods for our study, we conducted a literature search on Google Scholar with the query “((deep learning) AND ((computational histopathology) OR (whole slide images)) AND (classification))” for publications between 2017 and 2023. While recognizing the potential of hierarchical models to fuse local and global features, we focused on embedding aggregation technologies that do not require instance-level labels, drawing from studies that utilize embeddings at a uniform magnification for consistency. Our analysis centers on methods with public, implementable code and those frequently used in recent scholarly work for benchmarking purposes. Problem formulation of benchmarking aggregation methods is provided in Appendix A. Details on the selected aggregation methods and their computational complexities are provided in Appendix C. Table 2.
@@ 3 Summary of Datasets
1. BCa: Breast cancer (BCa) detection cohort. Breast cancer blocks and normal breast blocks were obtained from the pathology laboratory information system. A total of 1998 slides were sampled, with 999 positive and 999 negative. The positive slides were selected from blocks that received the routine biomarker panel for cancer cases (estrogen receptor: ER, progesterone receptor: PR, human epidermal growth factor receptor 2: HER2, and Ki67), while negative slides were selected from breast cases that did not have an order for the routine panel. Additionally, negative cases were selected if they were not mastectomy cases, did not have a synoptic report associated with the case, and had no mention of cancer or carcinoma in the report. 2. IBD: Inflammatory Bowel Disease (IBD) detection cohort. Normal mucosa samples were obtained from patients undergoing screening and routine surveillance lower endoscopy from 2018 to 2022. IBD cases, including first diagnoses and follow-ups, were included. A total of 1441 slides were sampled, 717 with active inflammation and 724 with normal mucosa. 3. BCa ER, BCa PR, BCa HER2: BCa biomarker prediction cohorts. Breast cancer cases are routinely assessed for ER, PR, and HER2 status using immunohistochemistry (IHC) and Fluorescence In Situ Hybridization (FISH). Results for each biomarker were automatically extracted from the pathology report. 4. BioMe BR HRD:: Breast (BR) Homologous Repair Deficiency (HRD) prediction cohort. Mount Sinai BioMe is a whole-exome sequencing cohort of 30k individuals, where carriers of pathogenic and protein-truncating variants affecting HRD genes, i.e., BRCA1 BRCA2 BRIP1 PALB2 RAD51 RAD51C RAD51D ATM ATR CHEK1 CHEK2, where included as positives. A subset of the BioMe dataset of patients with available breast pathology slides were included. Slides containing solely normal breast tissue and slides with breast cancer were both included. 5. LUAD EGFR: EGFR (Epidermal Growth Factor Receptor) mutation status prediction in Lung Adenocarcinoma (LUAD). Two datasets were collected, one from MSHS and one from MSKCC. For the MSHS dataset, a total of 294 slides were obtained from MSHS clinical slide database, 103 positive and 191 negative. The cohort was built following the guidelines described in previous work [31] to map mutations to a binary target. The MSKCC dataset consists of 1,000 slides with 307 positive and 693 negative and is a random subset of the dataset described in [31]. 6. NSCLC IO: Lung cancer immunotherapy response prediction. Non-small cell lung cancer (NSCLC) patients who received PD-L1 blockade-based immunotherapy were considered. Cytology specimens were excluded. The objective overall response was determined by RECIST [32] and performed by a blinded thoracic radiologist. A total of 454 slides were obtained, 86 positive and 368 negative.
@@ Acknowledgements
This work is supported in part through the use of research platform AI-Ready Mount Sinai (AIR.MS) and the expertise provided by the team at the Hasso Platner Institute for Digital Health at Mount Sinai (HPI.MS). This work was supported in part through the computational and data resources and staff expertise provided by Scientific Computing and Data at the Icahn School of Medicine at Mount Sinai and supported by the Clinical and Translational Science Awards (CTSA) grant UL1TR004419 from the National Center for Advancing Translational Sciences. Additionally, research funding was provided in part by a grant from from the National Library of Medicine (NLM) (R01 LM013766).